# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v1

@@ 3 Numerical solution
To solve Equation (20) numerically, we truncate and discretize the coordinate space from the real positive numbers to a finite grid of equally-spaced points $\{r_{n}\}_{n=0}^{M+1}$, where
@@
with $d$ a small discretization step. Our discretized coordinate space begins at the origin $r_{0}=0$, has $M$ points in its interior that we refer to as nodes, and ends at a finite but otherwise arbitrarily large radius $r_{M+1}=R$. We define the numerical reduced radial wavefunction in channel $i$ as $u_{i}(r_{n})$. Similarly, we define the numerical potential matrix element between the channels $i$ and $j$ as $V_{ij}(r_{n})$. We then use the Calculus of Finite Differences to approximate the second derivative of the wavefunction as
@@
To close the system of equations, we fix the values of the reduced radial wavefunction at the boundaries, $u_{i}(0)$ and $u_{i}(R)$. This is referred to as imposing Dirichlet boundary conditions. We can choose the appropriate boundary conditions using the known analytical behavior of the solutions to Equation (20) for $r\to 0$ and $r\to\infty$. Since a physical solution $u_{i}(r)$ goes to zero as $r^{l_{i}+1}$ for $r\to 0$, we impose
@@
The other boundary condition, $u_{i}(R)$, is more complicated. We define the threshold matrix as $T=\lim_{r\to\infty}V(r)$. For scattering problems,
$\lim_{r\to\infty}V_{ij}(r)=\delta_{ij}T_{i}$ (24)
so the threshold matrix $T$ is diagonal. Each threshold value $T_{i}$ is either a definite number or positive infinity. If $T_{i}$ is a finite number, then $\lvert V_{ii}(r)-T_{i}\rvert$ must approach zero faster than $1/r$ and $V_{ij}^{2}(r)/\lvert V_{jj}(r)-T_{i}\rvert$ must approach zero faster than $1/r$ for every $j\neq i$. The boundary condition $u_{i}(R)$ depends on whether the corresponding threshold $T_{i}$ is larger or smaller than the energy $E$ in the Schrödinger equation. If $T_{i}>E$, then a physical solution goes to zero for $r\to\infty$ and we impose $u_{i}(R)=0$. If otherwise $T_{i}\leq E$, a physical solution keeps oscillating indefinitely for $r\to\infty$. It is however a specific, finite number at the maximum distance $R$ in our truncated coordinate space. So, we impose $u_{i}(R)=b_{i}$, where $b_{i}$ is a real constant. We can conveniently impose the boundary condition for both $E\geq T_{i}$ and $E<T_{i}$ at once in the form
$u_{i}(R)=\theta(E-T_{i})b_{i},$ (25)
@@
$\theta(x)=\begin{cases}1&\text{if $x\geq 0$,}\\ 0&\text{if $x<0$,}\\ \end{cases}$ (26)
@@
Using Equations (22), (23), and (25), we derive from Equation (20) the numerical Scrödinger equation
$\sum_{j,m}\bigl{[}H_{ij,nm}-\delta_{ij}\delta_{nm}E\bigr{]}u_{j}(r_{m})\\ =\delta_{nM}\theta(E-T_{i})\frac{b_{i}}{2\mu_{i}d^{2}},$ (27)
@@
with $i,j=1,\dots,N$ denoting channels and $n,m=1,\dots,M$ denoting nodes. The Hamiltonian is a matrix with $N^{2}M^{2}$ elements, but only a small fraction of those are different from zero. The ratio of nonzero elements over the total scales like $1/M$, with $M$ being a very large number for a large maximum distance $R$ and a small discretization step $d$. The Hamiltonian in Equation (28) is therefore a sparse matrix.
@@
  • 1.
    The orbital angular momentum $l_{i}$, reduced mass $\mu_{i}$, and threshold $T_{i}$ for each channel $i$.
  • 2.
@@
The file channels.csv must have one line for the header followed by $N$ lines for the values. Each line must contain at least three entries. The header must contain at least the following strings: l, mu, and threshold. The following $N$ lines must contain at least the corresponding values for the channels $i=1,\dots,N$. An entry under l is an integer indicating $l_{i}$. An entry under mu is a decimal number indicating $\mu_{i}$. An entry under threshold is a decimal number indicating $T_{i}$ or inf if $T_{i}=\infty$. The ordering of the columns is irrelevant, but it must obviously be consistent between rows. Below is the typical structure of a minimal channels.csv file.
@@
SPARSE reads the values from the two CSV files and writes them inside two pandas DataFrame objects using the function read_csv from pandas [6]. From them, SPARSE calculates the Hamiltonian matrix elements $H_{ij,nm}$ from Equation (28) and writes them into a 2-dimensional NumPy array object [7] with elements $H_{\alpha\beta}$. This reshaping amounts to applying an invertible map $(i,n)\mapsto\alpha$ from pairs of indices $i=1,\dots,N$ and $n=1,\dots,M$ to a single index $\alpha=1,\dots,NM$. SPARSE reads the matrix elements from Equation (28) and writes them in the NumPy array following a C-style index order, which corresponds to setting
$\alpha=i+N(n-1).$ (29)
This choice of index order has the advantage that all the nonzero $H_{\alpha\beta}$ end up as close as possible to the main diagonal $\alpha=\beta$. From Equation (28), we see that the only nonzero $H_{ij,nm}$ are those with $\lvert i-j\rvert\leq N-1$ and $n=m$ or $\lvert n-m\rvert\leq 2$ and $i=j$. Then it follows from Equation (29) that the only nonzero $H_{\alpha\beta}$ are those with
$\lvert\alpha-\beta\rvert\leq\lvert i-j\rvert+N\lvert n-m\rvert\leq 2N.$ (30)
Therefore the $H_{\alpha\beta}$ form a banded matrix with $N$ nonzero upper diagonals and $N$ nonzero lower diagonals, $\alpha=\beta-\gamma$ and $\alpha=\beta+\gamma$ with $\gamma=1,\dots,N$.
SPARSE writes the Hamiltonian in the matrix diagonal ordered form, which is a 2-dimensional array consisting of only the $12N+$ 1 nonzero diagonals stacked from top to bottom. The upper and lower diagonals are padded so that they all have the same number of elements $NM$ as the main diagonal. The matrix diagonal ordered form has only $(2N+1)\times(NM)$ elements, and therefore it occupies significantly less memory than the full $(NM)\times(NM)$ matrix with its many zeroes. Consider for example a system with $N=4$ channels and $M=10^{6}$ nodes using 64-bit floating-point numbers. The matrix diagonal ordered form would take 288 megabytes of memory, opposed to a staggering 128 terabytes for the full matrix! Without the matrix diagonal ordered form, it would be all but impossible to accurately solve the Schrödinger equation using the finite-difference method.
So, we redefine the indices of Equation (27) as $(i,n)\mapsto\alpha$ and $(j,m)\mapsto\beta$ using Equation (29) and we obtain an ordinary nonhomogeneous linear system,
$\sum_{\beta}(H_{\alpha\beta}-\delta_{\alpha\beta}E)u_{\beta}=B_{\alpha}(E),$ (31)
@@
$B_{\alpha}(E)=\delta_{nM}\theta(E-T_{i})\frac{b_{i}}{2\mu_{i}d^{2}}.$ (32)
Note that if $E<T_{i}$ for all $i$, which corresponds to a bound state, then $B_{\alpha}(E)=0$ for all $\alpha$ and Equation (31) becomes an eigenvalue problem for the Hamiltonian matrix. Even though SPARSE is primarily designed for scattering problems, it also contains a handy bound-state calculator which solves Equation (31) with $B_{\alpha}(E)=0$ using the function eigsh from the sparse linear algebra module of SciPy [8].
[…]
@@ 4 Calculating the $K$-matrix
From now on we shall assume, without loss of generality, that the channels $i$ are ordered in increasing threshold value, $T_{i}\geq T_{j}$ for any $i\geq j$ [^2] This assumption is made in the paper for the sake of simplifying the notation. The SPARSE algorithm does not require the input channels to be sorted in any particular way. For a given energy $E$, the possible values of the constant in Equation (32) constitute a vector space of dimension $O$, where $O$ is the number of open channels $i$ whose threshold is less than $E$, $T_{i}\leq E$ for $i=1,\dots,O$. We can easily construct a basis for this vector space by substituting $b_{i}$ inside Equation (32) with
@@
where the first index $i=1,\dots,N$ labels the channels and the second index $j=1,\dots,O$ labels the vectors in the basis. We treat the $b_{ij}$ from Equation (33) as the elements of a $N\times O$ matrix $b$, and the result of plugging Equation (33) into (32) as the elements of a $(NM)\times O$ matrix $B$, dropping its dependence on $E$. We also define the $(NM)\times(NM)$ matrix
$A=H-EI.$ (34)
with $I$ the identity matrix. $A$ is a banded matrix which SPARSE calculates by shifting the main diagonal of the Hamiltonian by $-E$ in the matrix diagonal ordered form. With these definitions, we rewrite Equation (31) as a matrix equation
@@
For any energy $E$, SPARSE constructs the matrices $A$ and $B$ using Equations (32)-(34) and solves Equation (35) for the matrix $u$ using the function solve_banded from the linear algebra module of SciPy [8]. Then, SPARSE obtains the numerical wavefunctions from the matrix elements of $u$ using Equation (29), $u_{\alpha j}\to u_{ij}(r_{n})$. Next, we use these numerical solutions to calculate the reactance matrix $K$ for the scattering process. We will disregard from now on wavefunction components for confining channels with $T_{i}=\infty$, which are irrelevant to the scattering states at large $r$.
First, we define the numerical range of the potential $r_{V}$ as the maximum value of $r$ for which the potential matrix is significantly different from the threshold matrix (excluding diagonal elements for confining channels),
$r_{V}=\max\bigl{\{}r\;\big{|}\;\lVert V(r)-T\rVert>\epsilon\bigr{\}},$ (36)
@@
Next, we review the analytic asymptotic solution at large $r$. For $r>r_{V}$, Equation (20) reduces to a set of decoupled Schrödinger equations for free particles,
$\biggl{[}-\frac{1}{2\mu_{i}}\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}+\frac{l_{i}(l_{i}+1)}{2\mu r^{2}}+T_{i}-E\biggr{]}u_{i}(r)=0$ (37)
for every channel $i$ with a finite threshold $T_{i}$. Solutions to Equation (37) for closed channels $i=O+1,\dots,N$ with $T_{i}>E$ must approach zero at large distances and are irrelevant to the scattering states. For the open channels $i=1,\dots,O$ with $T_{i}\leq E$, Equation (37) has two linearly independent solutions. Using real boundary conditions and wavefunction normalization by energy (see Reference [9] for alternative conventions), the two solutions are
@@
where $p_{i}$ is the scattering momentum in the channel $i$,
$p_{i}=\sqrt{2\mu_{i}(E-T_{i})},$ (39)
@@
with $j_{l}(x)$ and $y_{l}(x)$ the spherical Bessel functions of the first and second kind. The asymptotic solution of any Schrödinger equation that coincides with Equation (37) at large $r$ can be written as a linear combination of the solutions in Equation (38). In particular, the asymptotic wavefunctions of a complete set of scattering states with energy $E$ can be expressed as
@@
Next, we compare the numerical solutions $u_{ij}(r_{n})$ to the analytic asymptotic solution in Equation (42) to extract the $K$-matrix elements. The numerical wavefunctions calculated from Equation (27) generally do not correspond to the asymptotic scattering states in Equation (42). Instead, the numerical wavefunctions at large $r$ correspond to a more general linear combination of Riccati-Bessel functions,
@@
The matrices $X$ and $Y$ depend on the matrix constant $B$ in Equation (35), but not the $K$-matrix. If instead of $B$ we had chosen any other admissible $B^{\prime}$, then instead of $X$ and $Y$ we would have obtained $X^{\prime}=XW$ and $Y^{\prime}=YW$ with $W$ an invertible $O\times O$ matrix. However, the $K$-matrix from Equation (44) would have been the same.
Finally, let us calculate explicitly the matrices $X$ and $Y$ from the numerical wavefunctions $u_{ij}(r_{n})$. Let us take the asymptotic solution in Equation (43) and take the further limit $x\gg l$ for the arguments of the Riccati-Bessel functions $S_{l}(x)$ and $C_{l}(x)$, which yields
@@
Then consider the integrals
$\int_{x_{0}}^{x_{0}+\pi}\sin^{2}(x)\,\mathrm{d}x=\int_{x_{0}}^{x_{0}+\pi}\cos^{2}(x)\,\mathrm{d}x=\frac{\pi}{2}$ (46)
and
$\int_{x_{0}}^{x_{0}+\pi}\sin(x)\cos(x)\,\mathrm{d}x=0,$ (47)
where $x_{0}$ can be any real number. Using Equations (46) and (47), we obtain
$X_{ij}$ $=\sqrt{\frac{2p_{i}^{3}}{\pi\mu_{i}}}\int_{a}^{b}u_{ij}(r)\sin\Bigl{(}p_{i}r-l_{i}\frac{\pi}{2}\Bigr{)}\,\mathrm{d}r,$ (48)
$Y_{ij}$ $=\sqrt{\frac{2p_{i}^{3}}{\pi\mu_{i}}}\int_{a}^{b}u_{ij}(r)\cos\Bigl{(}p_{i}r-l_{i}\frac{\pi}{2}\Bigr{)}\,\mathrm{d}r,$ (49)
where the limits of the integrals are
$a=R-\pi/p_{i}\qquad\text{and}\qquad b=R,$ (50)
with $R$ the maximum radius in the truncated coordinate space. Evaluating Equatons (48) and (49) using the numerical wavefunctions $u_{ij}(r_{n})$ and the trapezoidal rule provides a way of calculating $X_{ij}$ and $Y_{ij}$ that is both fast and numerically stable. SPARSE uses the trapezoid function from the integration module of SciPy [8].
5 Scattering amplitudes and resonances

### v2

@@ 3 Numerical solution
To solve Equation (20) numerically, we truncate and discretize the coordinate space for the distance $r$ from $\mathbb{R}_{0}^{+}$ to a finite grid of equally-spaced points $\{r_{n}\}_{n=0}^{M+1}$, where
@@
with $d$ a small discretization step. Our discretized coordinate space begins at the origin $r_{0}=0$, has $M$ points in its interior that we refer to as nodes, and ends at a finite but otherwise arbitrarily large radius $r_{M+1}=R$. The numerical reduced radial wavefunction in channel $i$ is $u_{i}(r_{n})$. The numerical potential matrix element between the channels $i$ and $j$ is $V_{ij}(r_{n})$. The numerical second derivative of the reduced radial wavefunction is evaluated using the leading-order finite-difference method,
@@
To close the system of equations, we fix the values of the reduced radial wavefunctions at the boundaries, $u_{i}(0)$ and $u_{i}(R)$. This is referred to as imposing Dirichlet boundary conditions. We can choose the physical boundary conditions using the known analytical behavior of the solutions to Equation (20) for $r\to 0$ and $r\to\infty$. Since a physical solution $u_{i}(r)$ goes to zero as $r^{l_{i}+1}$ for $r\to 0$, we impose
@@
The boundary condition at large $r$ is more complicated. For physical scattering problems, one has
$\lim_{r\to\infty}V_{ij}(r)=\delta_{ij}V_{i}^{\infty}$ (24)
where each $V_{i}^{\infty}$ is either a definite number or positive infinity. We identify $V_{i}^{\infty}$ with the energy threshold where the scattering channel $i$ opens up. The boundary condition for $u_{i}(R)$ depends on whether the energy $E$ in the Schrödinger equation is larger or smaller than the threshold $V_{i}^{\infty}$. If $E<V_{i}^{\infty}$, then a physical solution goes to zero for $r\to\infty$ and we impose $u_{i}(R)=0$. If otherwise $E>V_{i}^{\infty}$, a physical solution keeps oscillating indefinitely for $r\to\infty$. It is however a specific, finite number at the maximum distance $R$ in our truncated coordinate space. So, we impose $u_{i}(R)=b_{i}$, where $b_{i}$ is a real constant. We can conveniently impose the boundary condition for both cases $E<V_{i}^{\infty}$ and $E\geq V_{i}^{\infty}$ at once in the form
$u_{i}(R)=\theta(E-V_{i}^{\infty})b_{i},$ (25)
@@
[…]
@@
Using Equations (22), (23), and (25), we reduce Equation (20) down to a system of numerical Schrödinger equations,
$\sum_{j,m}\bigl[H_{ij,nm}-\delta_{ij}\delta_{nm}E\bigr]u_{j}(r_{m})\\ =\delta_{nM}\theta(E-V_{i}^{\infty})\frac{b_{i}}{2\mu_{i}d^{2}},$ (27)
@@
with $i,j=1,\dots,N$ denoting channels and $n,m=1,\dots,M$ denoting nodes. The numerical Hamiltonian is a matrix with $N^{2}M^{2}$ elements, but only a small fraction of them are different from zero. The ratio of nonzero elements over the total scales like $1/M$, with $M$ being a very large number for a large maximum distance $R$ and a small discretization step $d$. The numerical Hamiltonian in Equation (28) is therefore a sparse matrix.
@@
  • The orbital angular momentum $l_{i}$, reduced mass $\mu_{i}$, and threshold $V_{i}^{\infty}$ for each channel $i$.
@@
The file channels.csv must have one line for the header followed by $N$ lines for the values. Each line must contain at least three entries. The header must contain at least the following strings: l, mu, and threshold. The following $N$ lines must contain at least the corresponding values for the channels $i=1,\dots,N$. An entry under l is an integer indicating $l_{i}$. An entry under mu is a decimal number indicating $\mu_{i}$. An entry under threshold is a decimal number indicating $V_{i}^{\infty}$ or inf if $V_{i}^{\infty}=\infty$. The ordering of the columns is irrelevant, but it must obviously be consistent between rows. Below is the typical structure of a minimal channels.csv file.
@@
SPARSE reads the values from the two CSV files and stores them into two pandas DataFrame objects using the function read_csv from the pandas library [7]. From them, SPARSE calculates the Hamiltonian matrix elements $H_{ij,nm}$ from Equation (28) and writes them into a 2-dimensional NumPy array object [8] with elements $H_{\mu\nu}$. This reshaping amounts to applying an invertible map $(i,n)\mapsto\mu$ from pairs of indices $i=1,\dots,N$ and $n=1,\dots,M$ to a single index $\mu=1,\dots,NM$. SPARSE reads the matrix elements from Equation (28) and writes them in the NumPy array following a C-style index order, which corresponds to setting
$\mu=i+N(n-1).$ (29)
This choice of index order has the advantage that all the nonzero $H_{\mu\nu}$ end up as close as possible to the main diagonal. From Equation (28), we see that the only nonzero $H_{ij,nm}$ are those with $i=j$ and $\lvert n-m\rvert\leq 1$ or those with $n=m$ and $\lvert i-j\rvert\leq N-1$. Then it follows from Equation (29) that the only nonzero $H_{\mu\nu}$ are those with
$\lvert\mu-\nu\rvert\leq\lvert i-j\rvert+N\lvert n-m\rvert\leq N.$ (30)
Therefore the $H_{\mu\nu}$ form the elements of a banded matrix where all the nonzero elements are contained within the main diagonal, the first $N$ upper diagonals, and the first $N$ lower diagonals.
SPARSE writes the Hamiltonian in the matrix diagonal ordered form, which is a 2-dimensional array consisting of only the $2N+1$ nonzero diagonals stacked from top to bottom. The upper and lower diagonals are padded so that they all have the same number of elements $NM$ as the main diagonal. The matrix diagonal ordered form has only $(2N+1)\times(NM)$ elements, and therefore it occupies significantly less memory than the full $(NM)\times(NM)$ numerical Hamiltonian matrix with its many zeroes. Consider for example a system with $N=4$ channels and $M=10^{6}$ nodes using 64-bit floating-point numbers. The matrix diagonal ordered form would take 288 megabytes of memory, as opposed to a staggering 128 terabytes for the full matrix! If not for this memory optimization, it would be all but impossible to accurately solve a large system of Schrödinger equations using the finite-difference method.
We therefore redefine the indices of Equation (27) as $(i,n)\to\mu$ and $(j,m)\to\nu$ using Equation (29), which reshapes Equation (27) into an ordinary nonhomogeneous linear system,
$\sum_{\nu}(H_{\mu\nu}-\delta_{\mu\nu}E)u_{\nu}=B_{\mu}(E),$ (31)
@@
$B_{\mu}(E)=\delta_{nM}\theta(E-V_{i}^{\infty})\frac{b_{i}}{2\mu_{i}d^{2}}.$ (32)
Note that if $E<V_{i}^{\infty}$ for all $i$, which corresponds to a bound state, then $B_{\mu}(E)=0$ for all $\mu$ and Equation (31) becomes an eigenvalue problem for the Hamiltonian matrix. Even though SPARSE is primarily designed for calculating scattering states, it also contains a handy bound-state calculator which solves Equation (31) with $B_{\mu}(E)=0$ using the function eigsh from the sparse linear algebra module of SciPy [9].
Finally, a note about measuring units. SPARSE uses natural units, in which $\hbar=c=1$ and therefore $\text{Time}]=[\text{Length}]$, $\text{Mass}]=[\text{Energy}]$, and $\text{Length}]=[\text{Energy}]^{-1}$. There is only one independent dimension, typically chosen between $\text{Length}]$ in units of fm (femtometers) and $\text{Energy}]$ in units of eV (electron Volts). The conversion between length-based and energy-based units is easily achieved by means of the formula $\hbar c=1=197.327$ MeV fm. The SPARSE algorithm is agnostic to the user’s choice of units, but it assumes that the dimension of the inputs are $r_{n}]=[\text{Length}]$ and $V_{ij}(r_{n})]=[\mu_{i}]=[V_{i}^{\infty}]=[\text{Energy}]$.
@@ 4 Calculating the $K$-matrix
From now on we shall assume, without loss of generality, that the channels $i$ are ordered in increasing threshold value. It is worth noting that the SPARSE algorithm does not require the input channels to be sorted in any particular way. The assumption $V_{i}^{\infty}\geq V_{j}^{\infty}$ for any $i\geq j$ is only made to simplify the notation in this paper.
For a given energy $E$, the possible values of the constant in Equation (32) constitute a vector space of dimension $O$, where $O$ is the number of open channels $i$ whose threshold is less than $E$, $V_{i}^{\infty}\leq E$ for $i=1,\dots,O$. We can easily construct a basis for this vector space by substituting $b_{i}$ inside Equation (32) with
@@
where the first index $i=1,\dots,N$ labels the channels and the second index $j=1,\dots,O$ labels the vectors in the basis. Plugging Equation (33) into (32), we obtain a constant term with two indices, $B_{\mu j}(E)$, which we can treat as the elements of a $(NM)\times O$ matrix $B$, dropping its dependence on $E$. We also define the $(NM)\times(NM)$ matrix
A=HEI.A=H-EI. (34)
with $I$ the identity matrix. $A$ is a banded matrix which SPARSE calculates by shifting the main diagonal of the numerical Hamiltonian by E-E in the matrix diagonal ordered form. With these definitions, we rewrite Equation (31) as a matrix equation
@@
For any energy $E$, SPARSE constructs the matrices $A$ and $B$ using Equations (32)-(34) and solves Equation (35) for the matrix $u$ using the function solve_banded from the linear algebra module of SciPy [9]. Then, SPARSE reshapes the matrix $u$ into a set of numerical reduced wavefunctions using Equation (29): $\mu\to(i,n)$, $u_{\mu j}\to u_{ij}(r_{n})$. Next, we use these numerical wavefunctions to calculate the reactance matrix $K$ for the scattering process. We will disregard from now on wavefunction components for confining channels $i$ with $V_{i}^{\infty}=\infty$, which are irrelevant to the scattering states at large $r$.
First, we define the numerical range of the potential $r_{V}$ as the maximum value of $r$ for which any potential matrix element [^4] Excluding diagonal elements for confining channels. is significantly different from its limiting value at $r\to\infty$,
[…]
@@
Next, we review the analytic wavefunctions of the scattering states at large $r$. For $r\gg r_{V}$, Equation (20) reduces to a set of decoupled Schrödinger equations for free particles,
$\biggl[-\frac{1}{2\mu_{i}}\frac{\mathrm{d}^{2}\hphantom{r}}{\mathrm{d}r^{2}}+\frac{l_{i}(l_{i}+1)}{2\mu r^{2}}+V_{i}^{\infty}-E\biggr]u_{i}(r)=0,$ (37)
for every channel $i$ with a finite threshold $V_{i}^{\infty}$. For a closed channel $i$ with $V_{i}^{\infty}>E$ ($i=O+1,\dots,N$), a physical solution to Equation (37) must approach zero at large distances. For an open channel $i$ with $V_{i}^{\infty}\leq E$ ($i=1,\dots,O$), Equation (37) has two linearly independent physical solutions. Using real boundary conditions and wavefunction normalization by energy (see Reference [10] for alternative conventions), the two solutions are
@@
where $p_{i}$ is the scattering momentum in the open channel $i$,
$p_{i}=\sqrt{2\mu_{i}(E-V_{i}^{\infty})},$ (39)
@@
with $j_{l}(x)$ and $y_{l}(x)$ the spherical Bessel functions of the first and second kind, respectively. The asymptotic solution of any Schrödinger equation that coincides with Equation (37) at large $r$ can be written as a linear combination of the solutions in Equation (38). In particular, the asymptotic wavefunctions for a complete set of scattering states with energy $E$ can be expressed as
@@
Next, we compare the numerical wavefunctions $u_{ij}(r_{n})$ to the analytic asymptotic wavefunctions in Equation (42) to extract the $K$-matrix elements. The numerical wavefunctions calculated from Equation (35) generally do not correspond to the asymptotic scattering states in Equation (42). Instead, the numerical wavefunctions at large $r$ correspond to a more general linear combination of Riccati-Bessel functions,
@@
The matrices $X$ and $Y$ depend on the choice of the matrix constant $B$ in Equation (35), but the $K$-matrix does not. If instead of $B$ we had chosen any other admissible $B^{\prime}$, then instead of $X$ and $Y$ we would have obtained $X^{\prime}=XW$ and $Y^{\prime}=YW$ with $W=B^{-1}B^{\prime}$ an invertible $O\times O$ matrix. However, the $K$-matrix from Equation (44) would have been the same.
Finally, let us show how to calculate the matrices $X$ and $Y$ explicitly from the numerical wavefunctions $u_{ij}(r_{n})$. Consider the asymptotic wavefunction in Equation (43) and take the further limit $r\gg l_{i}/p_{i}$, which yields
@@
Now consider the following system of equations
$\left\{\begin{aligned} \mathcal{S}_{ij}&=\sqrt{\frac{2\pi p_{i}^{3}}{\mu_{i}}}\int_{\rho_{i}}^{R}u_{ij}(r)\sin\Bigl(p_{i}r-l_{i}\frac{\pi}{2}\Bigr)\,\mathrm{d}r\\ \mathcal{C}_{ij}&=\sqrt{\frac{2\pi p_{i}^{3}}{\mu_{i}}}\int_{\rho_{i}}^{R}u_{ij}(r)\cos\Bigl(p_{i}r-l_{i}\frac{\pi}{2}\Bigr)\,\mathrm{d}r\end{aligned}\right$., (46)
where the upper integration limit is the maximum radius $R$ and the lower integration limit $\rho_{i}$ satisfies
$\rho_{i}\gg r_{V}\qquad\text{and}\qquad\rho_{i}\gg l_{i}/p_{i},$ (47)
with $r_{V}$ the potential radius. The integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (46) are obviously convergent since the wavefunctions $u_{ij}(r)$ are linear combinations of sine and cosine functions in the integration interval. In fact, the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ can be exressed analytically in terms of $X_{ij}$ and $Y_{ij}$ by inserting Equation (45) and integrating the products of sine and cosine functions using
$\int_{x_{0}}^{x_{1}}\sin^{2}(x)\,\mathrm{d}x$ $=\frac{x_{1}-x_{0}}{2}-\frac{1}{4}\bigl(\sin(2x_{1})-\sin(2x_{0})\bigr),$ (48a)
$\int_{x_{0}}^{x_{1}}\cos^{2}(x)\,\mathrm{d}x$ $=\frac{x_{1}-x_{0}}{2}+\frac{1}{4}\bigl(\sin(2x_{1})-\sin(2x_{0})\bigr),$ (48b)
$\int_{x_{0}}^{x_{1}}\sin(x)\cos(x)\,\mathrm{d}x$ $=-\frac{1}{4}\bigl(\cos(2x_{1})-\cos(2x_{0})\bigr).$ (48c)
Thus Equation (46) becomes
$\left\{\begin{aligned} \mathcal{S}_{ij}&=(\alpha_{i}-\beta_{i})X_{ij}-\gamma_{i}Y_{ij}\\ \mathcal{C}_{ij}&=-\gamma_{i}X_{ij}+(\alpha_{i}+\beta_{i})Y_{ij}\end{aligned}\right$., (49)
where we have introduced the shorthand
$\alpha_{i}$ $=R-\rho_{i},$ (50a)
$\beta_{i}$ $=\frac{1}{2}\bigl(\sin(2R)-\sin(2\rho_{i})\bigr),$ (50b)
$\gamma_{i}$ $=\frac{1}{2}\bigl(\cos(2R)-\cos(2\rho_{i})\bigr).$ (50c)
The linear system in Equation (49) can be inverted to express $X_{ij}$ and $Y_{ij}$ in terms of the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$,
$\left\{\begin{aligned} X_{ij}&=\frac{1}{\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}}[(\alpha_{i}+\beta_{i})\mathcal{S}_{ij}+\gamma_{i}\mathcal{C}_{ij}]\\ Y_{ij}&=\frac{1}{\alpha_{i}^{2}-\beta_{i}^{2}-\gamma_{i}^{2}}[\gamma_{i}\mathcal{S}_{ij}+(\alpha_{i}-\beta_{i})\mathcal{C}_{ij}]\end{aligned}\right.$. (51)
Therefore, we can obtain the numerical values of $X_{ij}$ and $Y_{ij}$ by evaluating the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (46) using the numerical wavefunctions $u_{ij}(r_{n})$ and Simpson’s rule. SPARSE uses the simpson function from the integration module of SciPy [9]. As long as the momenta $p_{i}$ satisfy $p_{i}\ll d^{-1}$, the integrands of $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ are slowly oscillating functions with respect to the discretization step $d$. Then Simpson’s rule provides an accurate numerical estimate of the integrals $\mathcal{S}_{ij}$ and $\mathcal{C}_{ij}$ in Equation (46).
5 Scattering poles and amplitudes